Solve

x OVER 3 MINUS x OVER 4 = 3

Answer by checkley71(8403)

X/3-X/4=3
(4X-3X)/12=3 NOW CROSS MULTIPLY
4X-3X=3*12
X=36 ANSWER.
PROOF.
36/3-36/4=3
12-9=3
3=3